November 1996  ICTCM IX (Reno)  10

Explicit Solution — via Maple V (Release 3)

> SYS := { diff( x(t), t ) = v(t),
>  diff( v(t), t ) = -g - k/m*v(t) };

SYS := \{\frac{∂}{∂t} x( t) = v( t), \frac{∂}{∂t} v( t) = −g − \frac{k v( t)}{m}\}

> IC := { x(0)=x0, v(0)=0 };

IC := { x( 0 ) = x0 , v( 0 ) = 0 }

> K[pw_const] :=  k1+(k2-k1)*Heaviside(t-td);

K_{pw const} := k1 + ( k2 − k1 ) Heaviside( t − td )